# F-Actor: Controllable Conversational Behaviour in Full-Duplex Models

### v1

@@ A.1 Discrete Representations of Speech
The first category of discrete representations fo speech consists of units often termed “acoustic tokens” or Discrete Acoustic Units (DAUs), derived from reconstruction-based Neural Audio Codecs (NACs) such as Mimi, EnCodec, or SoundStream (Défossez et al., 2024, 2022; Zeghidour et al., 2021). These units are optimized to compress and perfectly reconstruct the input signal. Consequently, they encode all acoustic information, including speaker identity, prosody, and recording conditions, into a single stream that can be decoded directly back into a waveform. While this enables end-to-end modeling such as in Moshi and VALL-E (Défossez et al., 2024; Wang et al., 2023), the high information density of DAUs can make modeling stability challenging.
@@ A.2 Quantization Methods
The choice of quantization method fundamentally dictates the architecture of the Speech Language Model, specifically regarding how it models the multiple codes representing a single time step. The two primary paradigms involve dependent codebooks, typically resulting from Residual Vector Quantization (Zeghidour et al., 2021, RVQ), and independent codebooks, such as those derived from Finite Scalar Quantization (Mentzer et al., 2024, FSQ). In RVQ-based models like EnCodec or Mimi (Défossez et al., 2022, 2024), the codebooks operate sequentially where each quantizes the residual error of the previous one. This creates a strong dependency among the codes within a single frame. Therefore, generation often requires a hierarchical approach. The primary language model autoregressively predicts the code for the first RVQ layer, while a secondary mechanism, such as the Depth Transformer in Moshi or the nonautoregressive model in VALL-E, predicts the codes for subsequent RVQ layers for that frame (Défossez et al., 2024; Wang et al., 2023). Conversely, methods like NanoCodec use FSQ to project the latent space into independent subspaces, resulting in codebooks that are statistically independent (Casanova et al., 2025). This independence simplifies the modeling task by allowing the Language Model to predict all codes for a frame simultaneously. Such models, as seen in Hu et al. (2025), typically utilize a multiple head output layer, termed a Hydra head, which predicts the indices for all independent codebooks in parallel without the need for an additional sequential depthwise modeling step.
@@ C.2 Counting Backchannels and Interruptions
Parakeet provides the most reliable performance on Behavior-SD (Lee et al., 2025), missing on average only 1.0 backchannels and 0.5 interruptions per conversation. This means, 79.479.4% of the backchannels are correctly classified, and 79.279.2% of the interruptions are correctly classified. Results for all alignment methods using these threshold are reported in Tab.˜4. Based on these findings, we adopt Parakeet for all subsequent evaluations.
@@ Appendix D Training Details
Training and inference parameters are shown in Tab.˜5. We train our models on four NVIDIA A100-SXM4-40GB. Training takes around 46-48 hours, depending on the configuration.
For inference, we show that sampling parameters, in particular the temperature, have a great influence on model performance in Section˜7.3.1. We find that for our setup, a temperature of $0.9$ works best and hence report all results (except for those in Tab.˜6) with this temperature.
@@
*Unless explicitely stated otherwise.
@@ Appendix E Results
We report and discuss results for all of our models in Section˜7, with two exceptions, that are reported in the appendix: Detailed results using different temperatures during inference are shown in Tab.˜6. Results comparing models using NanoCodecs (FSQ) and Mimi (RVQ) audio encoding are shown in Tab.˜7.

### v2

@@ A.1 Discrete Representations of Speech
The first category of discrete representations fo speech consists of units often termed “acoustic tokens” or Discrete Acoustic Units (DAUs), derived from reconstruction-based Neural Audio Codecs (NACs) such as Mimi, EnCodec, or SoundStream (défossez2024moshispeechtextfoundationmodel; Défossez et al., 2022; Zeghidour et al., 2021). These units are optimized to compress and perfectly reconstruct the input signal. Consequently, they encode all acoustic information, including speaker identity, prosody, and recording conditions, into a single stream that can be decoded directly back into a waveform. While this enables end-to-end modeling such as in Moshi and VALL-E (défossez2024moshispeechtextfoundationmodel; Wang et al., 2023), the high information density of DAUs can make modeling stability challenging.
@@ A.2 Quantization Methods
The choice of quantization method fundamentally dictates the architecture of the Speech Language Model, specifically regarding how it models the multiple codes representing a single time step. The two primary paradigms involve dependent codebooks, typically resulting from Residual Vector Quantization (Zeghidour et al., 2021, RVQ), and independent codebooks, such as those derived from Finite Scalar Quantization (Mentzer et al., 2024, FSQ). In RVQ-based models like EnCodec or Mimi (Défossez et al., 2022; défossez2024moshispeechtextfoundationmodel), the codebooks operate sequentially where each quantizes the residual error of the previous one. This creates a strong dependency among the codes within a single frame. Therefore, generation often requires a hierarchical approach. The primary language model autoregressively predicts the code for the first RVQ layer, while a secondary mechanism, such as the Depth Transformer in Moshi or the nonautoregressive model in VALL-E, predicts the codes for subsequent RVQ layers for that frame (défossez2024moshispeechtextfoundationmodel; Wang et al., 2023). Conversely, methods like NanoCodec use FSQ to project the latent space into independent subspaces, resulting in codebooks that are statistically independent (Casanova et al., 2025). This independence simplifies the modeling task by allowing the Language Model to predict all codes for a frame simultaneously. Such models, as seen in Hu et al. (2025), typically utilize a multiple head output layer, termed a Hydra head, which predicts the indices for all independent codebooks in parallel without the need for an additional sequential depthwise modeling step.
@@ C.2 Counting Backchannels and Interruptions
Parakeet provides the most reliable performance on Behavior-SD (Lee et al., 2025), missing on average only 1.0 backchannels and 0.5 interruptions per conversation. This means, 79.479.4% of the backchannels are correctly classified, and 79.279.2% of the interruptions are correctly classified. Results for all alignment methods using these threshold are reported in Tab.˜5. Based on these findings, we adopt Parakeet for all subsequent evaluations.
@@ 3. Faithfulness: Does the dialogue avoid introducing contradictions or unrelated content?
C.3 Human Evaluation to Confirm the LLM Judge Findings
To confirm the reliability of the LLM judge, we conduct a human evaluation.
Setup. Annotators were given the same rating criteria as the LLM judge, evaluating each dialogue on its conversational coherence and faithfulness to the given narrative. We selected our two best-performing models (both using s/u loss, speaker embeddings, word alignment, and an audio delay of 2; one also incorporating BC and interruption tokens) alongside the Behaviour-SD original dialogues. Using the Pearmut platform (Zouhar and Kocmi, 2026), 12 annotators each rated 60 samples (20 dialogues × 3 models), with every dialogue annotated by 3 annotators, yielding 240 annotated instances across 80 unique dialogues.
Annotators. Annotators were 12 computer science researchers (5 female, 7 male), compensated according to the national collective wage agreement. Participants were informed that their ratings would be published, while their identities remain anonymous. No GDPR-protected personal data was collected. The annotators geographic location was in Europe.
Duration. The average annotation session (20 dialogues × 3 models) took 56.1 minutes per annotator, underscoring the practical necessity of automatic evaluation metrics for these kinds of evaluations.
The full annotation instructions and a platform screenshots can be found in Fig.˜5. The files used to set up the evaluation are also available in the github repository of this project.
Figure 5: Screenshot of the instructions and interface for the human evaluation.
@@ Appendix D Training Details
Training and inference parameters are shown in Tab.˜6. We train our models on four NVIDIA A100-SXM4-40GB. Training takes around 46-48 hours, depending on the configuration.
For inference, we show that sampling parameters, in particular the temperature, have a great influence on model performance in Section˜7.3.1. We find that for our setup, a temperature of $0.9$ works best and hence report all results (except for those in Tab.˜7) with this temperature.
@@
*Unless explicitly stated otherwise.
@@ Appendix E Results
We report and discuss results for all of our models in Section˜7, with two exceptions, that are reported in the appendix: Detailed results using different temperatures during inference are shown in Tab.˜7. Results comparing models using NanoCodecs (FSQ) and Mimi (RVQ) audio encoding are shown in Tab.˜8.